Which one is greater of the given two. 348  and 272. 
Please answer with steps

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
Take log and find value,
Y1 = 48 log 3 = 48* 0.4771 = 22.9
Y2 = 72 log 2 = 72 * 0.3010 = 21.672
Henec Y1 > Y2
Hence 3^48 is greater